OpenStudy (waheguru):

I dont understand this linear equation l=40-w

OpenStudy (anonymous):

What you are trying to do in this??

OpenStudy (anonymous):

If l and w are the length and width then: Sum of length and Width is 40..

OpenStudy (waheguru):

Where is the x

OpenStudy (waheguru):

For example y=mx+b THis would be y=40-b right but dont we need the x like the y

OpenStudy (ganpat):

Look at the graph...  linear equation need not be x.. it says different values of l at various w`s... for eg.. l = 40 - W 1. w= o, so l = 40... 2. w = 20, so l = 20 3. w = 40, so l = 0 the graph is linear...